OpenStudy (mchilds15):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form

OpenStudy (michele_laino):

If I consider a point P=(x,y) of your parabola, then applying the definition of a parabola, I can write this: \[(y-4)^{2}+(x+2)^{2}=(y-6)^{2}\] please develop this equality

OpenStudy (mchilds15):

So the vertex would be y=5 ..... and i think x=-2 Though standard form is \[f(x)=a(x-h)^2+k\] I still need (a) which i think is 1?

OpenStudy (mchilds15):

@campbell_st I need your help dude

OpenStudy (michele_laino):

from the equation above, I got: \[a=-\frac{ 1 }{ 4 }\]

OpenStudy (mchilds15):

... Are you sure?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

I can give you the equation of your parabola, namely: \[y=-\frac{ 1 }{ 4 }(x+2)^{2}+5\] is it that, please?

OpenStudy (mchilds15):

Why do you have two y values in the equation when I'm supposed to be using standard form? I'm confused how a is 1/4

OpenStudy (mchilds15):

I just don't understand the 1/4 out of all that

OpenStudy (michele_laino):

please note that I said that a=-1/4

OpenStudy (mchilds15):

Oh, sorry, a=-1/4 Could you explain to me how you got that answer?

OpenStudy (michele_laino):

please as I said before, please develop the above expression that I wrote in my first post

OpenStudy (mchilds15):

But how did you develop the expression?

OpenStudy (michele_laino):

for example: \[(y-4)^{2}=y ^{2}+16-8y\] now, please do the same with the remaining parentheses

OpenStudy (mchilds15):

... Oh